OpenStudy (anonymous):

use substitution to determine whether the point in on the line y=x+1; A(1,0)

OpenStudy (lgbasallote):

A(1,0) so substitute x = 1 and y = 0 y = x + 1 \[\implies 0 = 1 + 1\] is this true?

OpenStudy (anonymous):

No

OpenStudy (lgbasallote):

then the point is not on the line

OpenStudy (anonymous):

But I dont understand how to substitute to figure that out.

OpenStudy (lgbasallote):

A(1,0) is the point do you know which one is the x-coordinate and which one is the y coordinate?

OpenStudy (anonymous):

The X coordinate is 1? & the Y is 0? Sorry, im really bad at math

OpenStudy (lgbasallote):

but that is right

OpenStudy (lgbasallote):

so x coordinate is 1 and y coordinate is 0 your equation is y = x + 1 just substitute 1 into x and 0 into y that gives you 0 = 1 + 1 got it now?

OpenStudy (anonymous):

Yes, I understand. Thank you!

OpenStudy (lgbasallote):

welcome